# Parallel Split Learning with Global Sampling

### v1

@@ Abstract
The expansion of IoT devices and the demands of Deep Learning have highlighted significant challenges in Distributed Deep Learning (DDL) systems. Parallel Split Learning (PSL) has emerged as a promising derivative of Split Learning that is well suited for distributed learning on resource-constrained devices. However, PSL faces several obstacles, such as large effective batch sizes, non-IID data distributions, and the straggler effect. We view these issues as a sampling dilemma and propose to address them by orchestrating the mini-batch sampling process on the server side. We introduce the Uniform Global Sampling (UGS) method to decouple the effective batch size from the number of clients and reduce mini-batch deviation in non-IID settings. To address the straggler effect, we introduce the Latent Dirichlet Sampling (LDS) method, which generalizes UGS to balance the trade-off between batch deviation and training time. Our simulations reveal that our proposed methods enhance model accuracy by up to 34.1% in non-IID settings and reduce the training time in the presence of stragglers by up to 62%. In particular, LDS effectively mitigates the straggler effect without compromising model accuracy or adding significant computational overhead compared to UGS. Our results demonstrate the potential of our methods as a promising solution for DDL in real applications.
@@ I Introduction
Recent advances in Deep Learning (DL), driven by growing architecture sizes, have significantly impacted areas like audio signal processing, natural language processing, and smart healthcare [1, 2, 3, 4, 5]. This progress relies on large data volumes and substantial computing resources [6, 7, 8]. The expansion of Internet-of-Things (IoT) devices, such as smartphones and wearables, provides extensive data and computing capabilities, yet privacy laws often restrict access to this data [6]. Distributed Deep Learning (DDL) frameworks have emerged to address these challenges by enabling collaborative model training across multiple devices while preserving data privacy.
Federated Learning (FL) [9] is a prominent framework in DDL. In Federated Averaging (FedAvg), clients train model copies locally and periodically aggregate these models by averaging their parameters, facilitating global model convergence [9]. However, the high resource demands of FL restrict its implementation in edge computing and IoT environments, primarily due to substantial computational and communication burdens. Deploying FL on edge devices is further complicated by the straggler effect [10], which causes significant delays. Additionally, FL struggles with datasets that are small or non-independent and identically distributed (non-IID). Strategies such as client selection [11] and model quantization [12] help mitigate these issues, but do not resolve them completely.
Split Learning (SL) [13, 14] addresses some limitations of FL by partitioning neural network training between client devices and a central server. Clients conduct initial Forward Propagation (FP) on local data and send intermediate results to the server, which completes the FP, initiates the Backward Propagation (BP) and returns gradients to the clients to complete the BP. SL enhances data privacy, reduces computational demands on the client-side, and increases the convergence speed compared to FL, making it well-suited for edge and IoT applications [13, 15, 16]. However, it incurs high training and communication latency due to its sequential nature [15]. While SL performs better than FL with non-IID data [16, 17], it suffers from catastrophic forgetting as the number of clients increases [18].
To address training latency in SL, advanced derivatives of SL like Split Federated Learning (SFL) [19, 20, 21] and Parallel Split Learning (PSL) [22, 23, 24, 25, 26] have been proposed. SFL combines the techniques of SL and FL to facilitate parallel model training across multiple devices. This approach leverages the advantages of both FL and SL. However, it encounters increased communication overhead compared to SL because it requires both the transmission of intermediate model results from clients to the server and the aggregation of updated models back to clients. Furthermore, SFL necessitates that the server maintains individual model copies for each client, which complicates scalability. Although variants such as SFLv2, SFLv3 [20], and SFLG [21] aim to mitigate these drawbacks, practical challenges persist.
@@
PSL enhances parallelism without requiring multiple server model instances or client model aggregation. However, PSL still faces issues with large effective batch sizes, non-IID data, and the straggler effect:
  1. 1.
    Large Effective Batch Size Problem [27]: Also known as the Server-Client Update Imbalance Problem [25], where the effective batch size scales with the number of clients.
  2. 2.
    Non-Independent and Identically Distributed Data: Devices contribute data of varying distributions and sizes, causing batch distributions to deviate from the overall data distribution and leading to poor generalization.
  3. 3.
    Straggler Effect: Due to the need for synchronization, slower devices delay the training process, as the processing time of a single batch is determined by the slowest client, impacting throughput and total training time [10].
@@
We propose the Uniform Global Sampling (UGS) method as an extension of the PSL framework introduced in [22] to address the problems (1) and (2). UGS addresses problem (1) by allowing for variable local batch sizes that sum up to a fixed global size. This decouples the number of clients from the effective batch size. By analysing the deviation probability of global batches [^1] We use the term batch to refer to mini-batches in the context of DL. The difference between global and local batches is explained in Section III. from the overall data distribution, we show that the choice of the local batch sizes is crucial for the model convergence. In particular, we show that using fixed local batch sizes can increase the deviation probability of the global batches, especially in the presence of non-IID data. In UGS, the local batch sizes are sampled according to client selection probabilities. We show that if the client selection probabilities are proportional to the dataset sizes, the deviation bound of the global batches is at least as tight as for uniform sampling in central learning.
Inspired by the Latent Dirichlet Allocation model [28], we introduce the Latent Dirichlet Sampling (LDS) method as a generalization of UGS to address the straggler effect (3). LDS introduces a Dirichlet prior distribution over the client selection probabilities to control the incidence of stragglers in batches. The prior distribution is adjusted based on the delay times of the straggler clients and a trade-off hyperparameter. The client selection probabilities are estimated using the Expectation-Maximization (EM) algorithm for Maximum A Posteriori (MAP) estimation [29]. LDS balances the trade-off between batch deviation and training time by sampling local batch sizes according to these estimated client selection probabilities.
An overview of our methods is shown in Figure 1. Extensive simulation results demonstrate that the proposed UGS and LDS methods can significantly improve the accuracy and convergence speed of PSL, especially in the presence of non-IID data. Our results show that LDS effectively mitigates the straggler effect without compromising the model accuracy or adding significant computational overhead compared to UGS.
Contributions & Organization
@@
  • We analyse the deviation probability of global batches in PSL and show why the choice of the local batch size is crucial for the convergence of the model (Section IV-A).
@@
  • We introduce the UGS method to address the large effective batch size problem and the non-IID data problem in PSL (Section IV-B).
@@
  • Motivated by the straggler effect problem, we introduce LDS to balance the trade-off between batch deviation and training time, ensuring efficient and accurate model training (Section IV-C).
Our code and all the results of this work are publicly available on GitHub [30].
@@ II Related Work
PSL enables the parallel execution of the FP and BP on the clients and enhances scalability by avoiding the need for the server to maintain separate model copies for each client. Nevertheless, PSL faces the client decoupling problem [27] (also known as client detachment problem [25]), where the the information flow between clients is decoupled during the BP due to the independent execution of the BP on each client. Jeon and Kim [22] use gradient averaging over all clients to address the client decoupling problem. They chose the local batch sizes fixed and proportional to the dataset size of each client. Lyu et al. [31] propose averaging the clients’ local loss function outputs instead of their local gradients. This approach aligns with that of [22] due to the linearity of the gradient operator, but enhances privacy as clients can compute the local loss functions without sharing labels. [22] and [31] show consistent convergence and improved accuracy in both IID and non-IID settings, but do not address the large effective batch size problem.
@@
Pal et al. [27] propose SplitAvg to address the client decoupling problem. This method generalizes the approach in [22] by allowing averaging only a subset of the local gradients. Opposed to [22] and [31], the authors observe that using all clients for gradient averaging, the model may not converge. To address the large effective batch size problem, the authors propose scaling the learning rate on the server side while keeping the local batch size fixed and identical for all clients. Following the line of work from [26], Oh et al. [25] propose using smashed data augmentation, local loss function design and FedAvg for the client models to overcome the large effective batch size problem and client model decoupling.
@@
[27] and [25] provide promising solutions for the large effective batch size problem, but assume that the effective batch size scales with the number of clients. This assumption conflicts with the diminishing returns of larger batch sizes for gradient estimation [32] and the benefits of smaller batch sizes [33]. Furthermore, the empirical results are reported only for IID settings. They do not address the non-IID data problem, which is a significant challenge in PSL. For general SL, Cai and Wei [34] propose a method to handle non-IID data with a distillation loss function. The bargaining game approach for personalized SL [35] balances energy consumption and privacy in non-IID settings. However, these methods can not be directly applied to PSL.
@@
Other efforts to improve PSL focus on reducing training and communication latency [23, 24, 36]. Methods include last-layer gradient aggregation, resource optimization, client clustering and cut-layer selection. However, these methods do not address the straggler effect, which remains a significant challenge in PSL. A notable work in this direction is the BOA algorithm [37], which dynamically adjusts the batch size of slow clients in distributed learning systems to reduce the straggler effect.
@@
Different from other works, we approach the large effective batch size problem, the Non-IID data problem and the straggler effect as a sampling dilemma. By orchestrating the batch sampling process on the server side we allow variable local batch sizes and decouple the number of clients from the effective batch size. This enables global sampling methods to account for Non-IID data by reducing the deviation between the global batches and the overall data distribution and to mitigate the straggler effect by controlling the concentration of stragglers in global batches.
@@ III Preliminaries
We consider the PSL setting from [22] where a server and $K\in\mathbb{N}$ clients train a neural network model with parameter vector $\boldsymbol{w}$ collaboratively using the PSL framework. The set of clients is represented by $\mathcal{K}=\{1,2,\ldots,K\}$. We focus on supervised classification tasks, where the dataset $\mathcal{D}_{k}$ of each client $k\in\mathcal{K}$ consists of input-label pairs $\mathcal{D}_{k}=\{(\boldsymbol{x}_{k,i},y_{k,i})\}_{i=1}^{D_{k}}$, where $D_{k}\in\mathbb{N}$ is the dataset size, $\boldsymbol{x}_{k,i}$ is the input data, and $y_{k,i}\in\mathcal{M}$ is the ground truth class label of $\boldsymbol{x}_{k,i}$. Here, =$\mathcal{M}=\{1,2,\dots,M\}$ represents all classes and $M\in\mathbb{N}$ is the number of classes. The overall dataset $\mathcal{D}_{0}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$ is the union of all clients’ datasets, with a total size $D_{0}=\sum_{k=1}^{K}D_{k}$. We refer to the distribution of the datapoints in $\mathcal{D}_{0}$ as the empirical distribution. We define the class counting function $c:\mathbb{N}^{N}\times\mathbb{N}\to\mathbb{N}$ with $N\in\mathbb{N}$ as
$c(\boldsymbol{y},m)=\sum_{i=1}^{N}\delta(y_{i},m)$
@@
PSL mirrors SL by splitting the model parameters into a client-side lower segment $\boldsymbol{w}_{c}$ and a server-side upper segment $\boldsymbol{w}_{s}$ with $\boldsymbol{w}=(\boldsymbol{w}_{s},\boldsymbol{w}_{c})^{\top}$. The first layer on the server side is referred to as the cut layer. Before training, the server sends $\boldsymbol{w}_{c}$ to the clients, each storing a copy $\boldsymbol{w}_{c,k}$ with identical initial values. The function $f(\boldsymbol{x};\boldsymbol{w})$ maps $\boldsymbol{x}$ to a predicted value according to the given model parameters $\boldsymbol{w}$. Hence, the global model is given as $f(\boldsymbol{x};\boldsymbol{w})=f(f(\boldsymbol{x};\boldsymbol{w}_{c});\boldsymbol{w}_{s})$. When the argument of $f$ is a dataset or a subset of it, e.g. a batch $\mathcal{B}$ with batch size $B$, it operates element-wise on the input data and is defined as:
@@ IV-A Deviation Analysis
$P\left(\left|\frac{Y_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y_{m})}{B^{2}\epsilon^{2}}$
@@
$Y^{\prime}\sim\sum_{k=1}^{K}\text{Multinomial}(B_{k},\boldsymbol{\beta}_{k})$
@@
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y^{\prime}_{m})+(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}])^{2}}{B^{2}\epsilon^{2}}$
@@
If $B_{k}=B\frac{D_{k}}{D_{0}}$ or $\beta_{k,m}=\beta_{0,m}$ for all $k\in\mathcal{K}$ and $m\in\mathcal{M}$ then it holds that
@@ IV-B Uniform Global Sampling (UGS)
$\boldsymbol{z}^{(t)}=(z_{1}^{(t)},{z}_{2}^{(t)},\dots,{z}_{B}^{(t)})^{\top}$ are latent variables that represent the clients from which the $y_{i}^{(t)}$ are sampled. This allows the sampling process to be split between the server and the clients, where the server executes 2a) and the clients execute 2b). Since steps 2a) and 2b) are independent, we can postpone step 2b) and pre-compute the $B_{k}^{(t)}$ for all $k\in\mathcal{K}$ and $t\in\mathcal{T}$. Accordingly, we count the occurrences of each client $k$ in $\boldsymbol{z}^{(t)}$ with $B_{k}^{(t)}=c(\boldsymbol{z}^{(t)},k)$ right after step 2a). The server then sends the $B_{k}^{(t)}$ to the clients.
Once training starts, the clients can execute the step 2b) on demand, i.e., sampling local batches of size $B_{k}^{(t)}$. This equals substituting $B_{k}$ with $B_{k}^{(t)}$ in the first step of PSL in Section III. The rest of the training process remains unchanged. Consequently, the effective batch size $B$ becomes a hyperparameter that is no longer dependent on the number of clients or their respective dataset sizes, addressing the large effective batch size problem.
@@
$\mathbb{P}\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y_{m})}{B^{2}\epsilon^{2}}$

### v2

@@ Abstract
The expansion of IoT devices and the demands of deep learning have highlighted significant challenges in distributed deep learning systems. Parallel split learning has emerged as a promising derivative of split learning well suited for distributed learning on resource-constrained devices. However, parallel split learning faces several challenges, such as large effective batch sizes, non-independent and identically distributed data, and the straggler effect. We view these issues as a sampling dilemma and propose to address them by orchestrating a mini-batch sampling process on the server side. We introduce a new method called uniform global sampling to decouple the effective batch size from the number of clients and reduce the mini-batch deviation. To address the straggler effect, we introduce a novel method called Latent Dirichlet Sampling, which generalizes uniform global sampling to balance the trade-off between batch deviation and training time. Our simulations reveal that our proposed methods enhance model accuracy by up to 34.1% in non-independent and identically distributed settings and reduce the training time in the presence of stragglers by up to 62%. In particular, Latent Dirichlet Sampling effectively mitigates the straggler effect without compromising model accuracy or adding significant computational overhead compared to uniform global sampling. Our results demonstrate the potential of our methods to mitigate common challenges in parallel split learning.
@@ I Introduction
In recent years, deep learning (DL) has undergone transformative growth in various domains [1, 2, 3, 4, 5]. Distributed deep learning (DDL) has emerged as a powerful paradigm in the era of the Internet of Things (IoT) and edge computing, addressing the challenges posed by the proliferation of resource-constrained devices and the exponential growth of data at the network edge [6, 7, 8]. This approach enables the distribution of computationally intensive DL tasks across multiple nodes, including IoT devices and edge servers, to overcome the limitations of centralized learning (CL). Using the collective computational power of interconnected devices, DDL not only enhances the efficiency of model training and inference, but also reduces latency, improves privacy, and optimizes resource utilization.
Federated learning (FL) [9] is a prominent framework in DDL. In federated averaging (FedAvg), clients train model copies locally and periodically aggregate these models by averaging their parameters, facilitating global model convergence [9]. However, the high resource demands of FL restrict its implementation in edge computing and IoT environments, primarily due to substantial computational and communication burdens. Strategies such as client selection [10] and model quantization [11] help mitigate these issues but do not resolve them completely.
Split learning (SL) [12, 13] addresses some limitations of FL by partitioning neural network training between client devices and a central server. Clients conduct initial forward propagation (FP) on local data and send intermediate results to the server, which completes the FP, initiates the backward propagation (BP), and returns gradients to the clients to complete the BP. SL enhances data privacy, reduces computational demands on the client side, and increases the convergence speed compared to FL, making it well-suited for edge and IoT applications [12, 14, 15]. However, it incurs high training and communication latency due to its sequential nature [14].
To address training latency in SL, advanced derivatives of SL like split federated learning (SFL) [16, 17, 18] and parallel split learning (PSL) [19, 20, 21, 22, 23] have been proposed. SFL combines the SL and FL techniques to facilitate parallel model training across multiple devices. However, SFL requires that the server maintains individual model copies for each client and the aggregation of updated client models, which complicates scalability.
@@
PSL enhances parallelism without the need for multiple server model instances or client model aggregation. However, PSL still encounters challenges, including the large effective batch size problem [24], also known as the server-client update imbalance problem [22], where the effective batch size scales with the number of clients. Additionally, devices contribute data of varying distributions and sizes, causing batch distributions to deviate from the overall data distribution, which can lead to poor generalization. This issue of non-IID data is a general problem in DDL, not specific to PSL. Another challenge is the straggler effect [25], where slower devices delay the training process due to the need for synchronization, as the processing time of a single batch is determined by the slowest client, impacting throughput and total training time.
@@
To address the large effective batch size problem and the issue of non-IID data, we propose a new method called uniform global sampling (UGS) as an extension to the PSL framework introduced in [19]. UGS allows for variable local batch sizes that sum up to a fixed global batch size. This decouples the number of clients from the effective batch size. By analysing the deviation probability of global batches [^1] We use the term batch to refer to mini-batches in the context of DL. The difference between global and local batches is explained in Section III. from the overall data distribution, we show that the choice of the local batch sizes is crucial for model convergence. In particular, we show that using fixed local batch sizes can increase the deviation probability of the global batches, especially in the presence of non-IID data. In UGS, the local batch sizes are sampled according to client selection probabilities. We show that if the client selection probabilities are proportional to the dataset sizes, the deviation bound of the global batches is at least as tight as for uniform sampling in central learning. Inspired by the Latent Dirichlet Allocation model [26], we introduce a second novel method named Latent Dirichlet Sampling (LDS) that generalizes UGS to address the straggler effect. LDS introduces a Dirichlet prior distribution over the client selection probabilities to control the concentration of stragglers in global batches. The prior distribution is adjusted based on the delay times of the straggler clients and a trade-off hyperparameter. The client selection probabilities are estimated using the expectation-maximization (EM) algorithm for maximum a posteriori (MAP) estimation [27]. LDS balances the trade-off between batch deviation and training time by sampling local batch sizes according to these estimated client selection probabilities.
An overview of our methods is shown in Figure 1. The main contributions of this paper are listed below:
@@
  • We analyse the deviation probability of global batches in PSL and show why the choice of the local batch size is crucial for the convergence of the model.
@@
  • We introduce the UGS method to address the large effective batch size problem and the non-IID data problem in PSL.
@@
  • Motivated by the straggler effect problem, we introduce LDS to balance the trade-off between batch deviation and training time, ensuring efficient and accurate model training.
  • Our simulation results show that the UGS and LDS methods significantly improve PSL accuracy and convergence speed, especially with non-IID data. LDS effectively mitigates the straggler effect without compromising model accuracy or adding significant computational overhead compared to UGS.
The remainder of this paper is organized as follows. Sections II and III present related works and preliminaries of our work. In Section IV, our novel UGS and LDS methods are introduced after analyzing the mini-batch deviation in PSL. Section V presents the simulation results and Section VI concludes the paper with final remarks.
@@ II Related Work
Initial versions of PSL faced the client decoupling problem [24] (also known as client detachment problem [22]), where the information flow between clients is decoupled during the BP due to independent execution. Jeon and Kim [19] use gradient averaging over all clients to address the client decoupling problem. They chose the local batch sizes fixed and proportional to the dataset size of each client. Lyu et al. [28] propose averaging the clients’ local loss function outputs instead of their local gradients. This approach aligns with that of [19] due to the linearity of the gradient operator, but enhances privacy as clients can compute the local loss functions without sharing labels. Both show consistent convergence and improved accuracy in both IID and non-IID settings but do not address the large effective batch size problem.
@@
Pal et al. [24] propose gradient averaging over a subset of the clients. Opposed to [19] and [28], the authors observe that the model may not converge if gradients are averaged over all clients. To address the large effective batch size problem, the authors propose scaling the learning rate on the server side while keeping the local batch size fixed and identical for all clients. Following the line of work from [23], Oh et al. [22] propose using smashed data augmentation, local loss function design and FedAvg for the client models to overcome the large effective batch size problem and client model decoupling.
@@
Pal et al. [24] and Oh et al. [22] provide promising solutions for the large effective batch size problem, but assume that the effective batch size scales with the number of clients. However, it has been shown that increasing the batch size leads to diminishing returns for gradient estimation [29] and that smaller batch sizes [30] can be more beneficial for model generalization. Furthermore, the empirical results are reported only for IID settings. They do not address the non-IID data problem, which is a significant challenge in PSL.
Cai and Wei [31] propose an incentive mechanism to handle non-IID data in PSL, which encourages clients with larger and more evenly distributed datasets to participate in training. The bargaining game approach for personalized SL [32] balances energy consumption and privacy in non-IID settings. However, these methods may face scalability issues and may not fully address the heterogeneity of data distributions across diverse clients.
@@
Other efforts to improve PSL focus on reducing training and communication latency [20, 21, 33]. Methods include last-layer gradient aggregation, resource optimization, client clustering, and cut-layer selection. However, these methods do not address the straggler effect, which remains a significant challenge in PSL. A notable related work is the BOA algorithm [34], which dynamically adjusts the batch size of slow clients in distributed learning systems to reduce the straggler effect.
@@
Contrary to other works, we view the large effective batch size problem, the non-IID data problem, and the straggler effect as a sampling dilemma. By orchestrating the batch sampling process on the server side, we allow variable local batch sizes and decouple the number of clients from the effective batch size. This enables global sampling methods to account for non-IID data by reducing the deviation between the global batches and the overall data distribution and mitigating the straggler effect by controlling the concentration of stragglers in global batches.
@@ III Preliminaries
We consider the PSL setting from [19] where a server and $K\in\mathbb{N}$ clients train a neural network model with parameter vector $\boldsymbol{w}$ collaboratively using the PSL framework. The set of clients is represented by $\mathcal{K}=\{1,2,\ldots,K\}$. We focus on supervised classification tasks, where the dataset $\mathcal{D}_{k}$ of each client $k\in\mathcal{K}$ consists of input-label pairs $\mathcal{D}_{k}=\{(\boldsymbol{x}_{k,i},y_{k,i})\}_{i=1}^{D_{k}}$, where $D_{k}\in\mathbb{N}$ is the dataset size, $\boldsymbol{x}_{k,i}$ is the input data, and $y_{k,i}\in\mathcal{M}$ is the ground truth class label of $\boldsymbol{x}_{k,i}$. Here, =$\mathcal{M}=\{1,2,\dots,M\}$ represents all classes and $M\in\mathbb{N}$ is the number of classes. The overall dataset $\mathcal{D}_{0}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$ is the union of all clients’ datasets, with a total size $D_{0}=\sum_{k=1}^{K}D_{k}$. We refer to the distribution of the data points in $\mathcal{D}_{0}$ as the empirical distribution. We define the class counting function $c:\mathbb{N}^{d}\times\mathbb{N}\to\mathbb{N}$ with $d\in\mathbb{N}$ as
$c(\boldsymbol{y},m)=\sum_{i=1}^{d}\delta(y_{i},m)$
@@
PSL mirrors SL by splitting the model parameters into a client-side lower segment $\boldsymbol{w}_{c}$ and a server-side upper segment $\boldsymbol{w}_{s}$ with $\boldsymbol{w}=(\boldsymbol{w}_{s},\boldsymbol{w}_{c})^{\top}$. The first layer on the server side is referred to as the cut layer. Before training, the server sends $\boldsymbol{w}_{c}$ to the clients, each storing a copy $\boldsymbol{w}_{c,k}$ with identical initial values. The function $f(\boldsymbol{x};\boldsymbol{w})$ maps $\boldsymbol{x}$ to a predicted value according to the given model parameters $\boldsymbol{w}$. Hence, the global model is given as $f(\boldsymbol{x};\boldsymbol{w})=f(f(\boldsymbol{x};\boldsymbol{w}_{c});\boldsymbol{w}_{s})$. When the argument of $f$ is a dataset or a subset of it, e.g., a batch $\mathcal{B}$ with batch size $B$, it operates element-wise on the input data and is defined as:
@@ IV-A Deviation Analysis
$P\left(\left|\frac{Y_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y_{m})}{B^{2}\epsilon^{2}}.$
@@
$Y^{\prime}\sim\sum_{k=1}^{K}\text{Multinomial}(B_{k},\boldsymbol{\beta}_{k}).$
@@
$P\left(\left|\frac{Y^{\prime}_{m}}{B}-\beta_{0,m}\right|\geq\epsilon\right)\leq\frac{\textnormal{Var}(Y^{\prime}_{m})+(\mathbb{E}[Y^{\prime}_{m}]-\mathbb{E}[Y_{m}])^{2}}{B^{2}\epsilon^{2}}.$
@@
If $B_{k}=B\frac{D_{k}}{D_{0}}$ or $\beta_{k,m}=\beta_{0,m}$ for all $k\in\mathcal{K}$ and $m\in\mathcal{M}$, then it holds that
@@ IV-B Uniform Global Sampling (UGS)
$\boldsymbol{z}^{(t)}=(z_{1}^{(t)},{z}_{2}^{(t)},\dots,{z}_{B}^{(t)})^{\top}$ represent the clients from which the $y_{i}^{(t)}$ are sampled. This allows the sampling process to be split between the server and the clients, where the server executes 2a) and the clients execute 2b). Since steps 2a) and 2b) are independent, we can postpone step 2b) and pre-compute the $B_{k}^{(t)}$ for all $k\in\mathcal{K}$ and $t\in\mathcal{T}$. Accordingly, we count the occurrences of each client $k$ in $\boldsymbol{z}^{(t)}$ with $B_{k}^{(t)}=c(\boldsymbol{z}^{(t)},k)$ right after step 2a). The server then sends the $B_{k}^{(t)}$ to the clients.
Once training starts, the clients can execute step 2b) on demand, i.e., sampling local batches of size $B_{k}^{(t)}$. This equals substituting $B_{k}$ with $B_{k}^{(t)}$ in the first step of PSL in Section III. The rest of the training process remains unchanged. Consequently, the effective batch size $B$ becomes a hyperparameter that is no longer dependent on the number of clients or their respective dataset sizes, addressing the large effective batch size problem.